1. On January 1, $808,000, five-year, 10% bonds, were issued for $783,760. Interest is paid semiannually on January 1 and July 1. If the issuing corporation uses the straight-line method to amortize the discount on bonds payable, the semiannual amortization amount is

a.$40,400

b.$4,848

c.$24,240

d.$2,424

2. Bonds Payable has a balance of $991,000 and Premium on Bonds Payable has a balance of $10,901. If the issuing corporation redeems the bonds at 103, what is the amount of gain or loss on redemption?

a.$18,829 loss

b.$10,901 loss

c.$1,020,730 gain

d.$10,901 gain
Answer #1

Answers

--Requirement 1 and 2, with calculation and solution

A Face value $808,000
B Issue Price $783,760
C = A - B Discount on issue $24,240
D No of years                                    5
E = Dx 2 No of semi annual payments                                 10
F = C/E Straight lime semi annual amortisation $2,424
Correct Answer Option 'D' $ 2424
A Bonds Payable face value $991,000
B Unamortised Premium $10,901
C = A+B Total Carrying value of Liability $1,001,901
D = A x 103/100 Cash paid to pay off liability $1,020,730
E = D - C Loss on redemption $18,829
Correct Answer Option 'A' $ 18829 Loss